### v1

Abstract
The (strong) isometric path complexity is a recently introduced graph invariant that captures how arbitrary isometric paths (i.e. shortest paths) of a graph can be viewed as a union of few “rooted" isometric paths (i.e. isometric paths with a common end-vertex). It is known that this parameter can be computed optimally in polynomial time. Seemingly unrelated graph classes studied in metric graph theory (e.g. hyperbolic graphs), geometric intersection graph theory (e.g. outerstring graphs), and structural graph theory (e.g. (theta, prism, pyramid)-free graphs) have been shown to have bounded strong isometric path complexity [Chakraborty et al., MFCS ’23].
We show that important graph classes studied in coarse graph theory (as introduced by [Georgakopoulos & Papasoglu ’23]) have bounded strong isometric path complexity. Specifically, we show that the strong isometric path complexity of $K_{2,t}$-asymptotic minor-free graphs is bounded. In fact, we prove a more general result. Let $U_{t}$ denote the graph obtained by adding a universal vertex to a path of $t-1$ edges. We show that the strong isometric path complexity of $U_{t}$-asymptotic minor-free graphs is bounded. This implies that $K_{4}^{-}$-asymptotic minor-free graphs, i.e. graphs that are quasi-isometric to a cactus [Fujiwara & Papasoglu ’23] have bounded strong isometric path complexity. On the other hand, $K_{4}$-minor-free graphs have unbounded strong isometric path complexity.
We also show that graphs whose all induced cycles of length at least 4 have the same length (also known as monoholed graphs as defined by [Cook et al., JCTB ’24]) form a subclass of $U_{4}$-asymptotic minor-free graphs. Hence, the strong isometric path complexity of monoholed graphs is bounded. On the other hand, we show that even-hole free graphs have unbounded strong isometric path complexity.
We also investigate which graph operations preserve strong isometric path complexity. We show that the strong isometric path complexity is preserved under the fixed power and line graph operators, two important graph operations. We also show that the clique-sums of finitely many graphs with strong isometric path complexity at most $k$, yields a graph with strong isometric path complexity at most $183k+$ 18.
1 Introduction
The study of the distance-metric of graphs is at the heart of metric graph theory [7] and finds numerous practical applications. To captures how arbitrary isometric paths (i.e. shortest paths) of an unweighted graph can be viewed as a union of few “rooted" isometric paths (i.e. isometric paths with a common end-vertex), a graph invariant known as isometric path complexity was introduced in [10]. Formally, it is defined as follows.
Given an unweighted and undirected graph $G$ and a vertex $v\in V(G)$, an isometric path $P$ of $G$ is $v$-rooted if $v$ is one of the end-vertices of $P$. The isometric path complexity of a connected graph $G$, denoted by $ipco\left(G\right)$, is the minimum integer $k$ such that there exists a vertex $v\in V(G)$ satisfying the following property: the vertices of any isometric path of $G$ can be covered by $k$ many $v$-rooted isometric paths.
A stronger variant of the isometric path complexity which is the main topic of this paper is also defined in [10] as follows.
Definition 1.
The strong isometric path complexity of a connected graph $G$, denoted by $sipco\left(G\right)$, is the minimum integer $k$ such that every vertex $v\in V(G)$ satisfies the following property: the vertices of any isometric path of $G$ can be covered by $k$ many $v$-rooted isometric paths.
Clearly, for any graph $G$, $ipco\left(G\right)\leq sipco\left(G\right)$. We say that the isometric path complexity (resp. strong isometric path complexity) of a graph class $\mathcal{G}$ is bounded if there is a constant $k$ that depends only on $\mathcal{G}$ such that for all $G\in\mathcal{G}$, the isometric path complexity (resp. strong isometric path complexity) is at most $k$.
Motivation: Originally, these parameters were conceived to propose approximation algorithms for the problem Isometric Path Cover, (also known as Shortest Path Cover) where the objective is to find a minimum-cardinality set of isometric paths so that each vertex of the graph belongs to at least one of the solution paths. Isometric Path Cover has been studied recently in the algorithmic graph theory community [11, 10, 12] and finds applications in machine learning [32], cop decompositions (in relation with the celebrated Cops and Robber game [1]), bus route design [11] etc. It was shown that Isometric Path Cover admit a polynomial-time $c$-approximation algorithm on graphs with isometric path complexity at most $c$ [10] (while being NP-hard even on chordal graphs). Unlike some other metric graph parameters (e.g. treelength), the (strong) isometric path complexity can be computed optimally in polynomial time [10].
From the structural perspective, it turns out that many important graph classes such as hyperbolic graphs, outerstring graphs, and (theta,prism,pyramid)-free graphs have bounded strong isometric path complexity [10]. These graph classes are incomparable to each other with respect to inclusion, and isometric path complexity being bounded is perhaps the first known interesting property shared by these seemingly unrelated graph classes. Hyperbolic graphs are important in metric graph theory [7, 26] and include classic graph classes such as chordal graphs, and more generally, graphs with bounded tree-length). But hyperbolic graphs exclude arbitrarily large cycles as isometric subgraphs. Outerstring graphs (and its subclasses) are studied in geometric intersection graph theory and computational geometry [6, 8, 9, 31]. The class of (theta,prism,pyramid)-free graphs has been studied in structural graph theory [2, 15]. Note that (as opposed to hyperbolic graphs) outerstring graphs and (theta,prism,pyramid)-free graphs may contain arbitrarily many large isometric cycles, although in a structured manner. See Figure 1 for graph classes with bounded strong isometric path complexity. In this paper, we show that the strong isometric path complexity is also important in coarse graph theory.
Bounded strong isometric path complexity
$U_{t}$-asymptotic minor-free *
Bounded hyperbolicity [10]
$K_{4}$-minor free [10]
Perfect
Bipartite
Even-hole free *
($\ell$-theta, $\ell$-prism, $\ell$-pyramid)-free [10]
Outerstring [10]
Outerplanar [10]
Bounded tree-length [11]Bounded chordality [11]Bounded diameter [11]
Monoholed *
Chordal [11]
Line graph *
Graph powers *
Clique-sum *
Figure 1: Inclusion diagram for graph classes. If a class $A$ has an upward path to class $B$, then $A$ is included in $B$. Constant bounds for the strong isometric path complexity on graph classes and graph operations marked with * are contributions of this paper. The boxes marked in gray indicate graph classes with unbounded isometric path complexity. Elliptical boxes represent graph operations.
Asymptotic minors: Recently, Georgakopoulos and Papasoglu [25] gave an overview of the area called coarse graph theory, where the objective is to “look at graphs from far away” and to understand their “large-scale geometry” [25]. Many problems and questions posed in [25] involve asymptotic minors and fat minors, and have garnered recent attention [24, 3, 5, 4, 20]. A graph $G$ contains a graph $H$ as a $K$-fat minor if it is possible to find in $G$, (i) a collection of connected subgraphs called branch sets corresponding to $V(H)$, and a collection of connected sets called branch paths corresponding to $E(H)$ such that (i) after contracting each branch set to a vertex, and each branch path to an edge, we obtain a copy of $H$, and (ii) the above sets are at pairwise distance at least $K$, except for incident branch set–branch paths pairs, which must be at distance 00 by definition. A graph class $\cal G$ contains $H$ as an asymptotic minor if for every integer $K$, there exists a graph $G\in\cal G$, such that $G$ contains $H$ as a $K$-fat minor. Otherwise, $\cal G$ is $H$-asymptotic minor-free.
It was conjectured in [25] that $H$-asymptotic minor-free graphs are quasi-isometric to $H$-minor-free graphs for every graph $H$. (A quasi-isometric mapping of a graph $G$ to a graph $H$ is a mapping of $V(G)$ to $V(H)$ that preserves the distance up to a constant multiplicative and additive distortion.) If true, coarse analogues of various results from graph minor theory would be true as well. However, the above conjecture has been proven to be false in general [20]. On the other hand, the above conjecture is true for $H=K_{4}$ [4], $H=K_{4}^{-}$ [24, 4], and $H=K_{2,3}$ [14].
In this paper, we consider a super class of $K_{2,3}$-asymptotic minor-free graphs and $K_{4}^{-}$-asymptotic minor free graphs and show that its strong isometric path complexity is bounded. Let $U_{t}$ denote the graph obtained by taking a path of length $t$ and introducing a universal vertex. See Figure 2 for an example. We prove the following.
Figure 2: The graph $U_{5}$.
Theorem 1.
There is a function $f\colon\mathbb{N}\times\mathbb{N}\rightarrow\mathbb{N}$ such that the strong isometric path complexity of graphs that do not contain $U_{t}$ as a $K$-fat minor is at most $f(K,t)$.
We note that $K_{2,t}$-asymptotic minor-free graphs are also $U_{t}$-asymptotic minor-free graphs. (See Proposition 1 for a proof.) Hence, Theorem 1 implies the following.
Corollary 1.
There is a function $f\colon\mathbb{N}\times\mathbb{N}\rightarrow\mathbb{N}$ such that the strong isometric path complexity of graphs that do not contain $K_{2,t}$ as a $K$-fat minor is at most $f(K,t)$.
The class of $K_{2,3}$-asymptotic minor-free graphs and its subclasses (e.g. $K_{2,3}$-induced minor-free graphs, universally-signable graphs, outerstring graphs) have been studied in the literature [14, 9, 6, 17, 29, 19]. We note that $K_{2,t}$-asymptotic minor-free graphs are strictly included in the class of $U_{t}$-asymptotic minor-free graphs. For an integer $k\geq 2$, let $F_{k}$ denote the graph obtained by taking $k$ copies of paths of length $k$ and identifying the end-vertices of the paths. Let $\mathcal{F}=\displaystyle\bigcup\limits_{k\geq 2}F_{k}$. Clearly, $\cal F$ is $U_{t}$-asymptotic minor-free but contains $K_{2,t}$ as an asymptotic minor for all $t\geq 1$. observe that all graphs in $\cal F$ are monoholed i.e. the lengths of all induced cycles in the graph are the same.
Monoholed graphs and even-hole free graphs: A hole is an induced cycle of length at least 4. The study of graphs whose hole lengths are restricted has been pursued in Structural Graph Theory [18, 27, 35]. A graph $G$ is $\ell$-holed if all holes of $G$ are of length $\ell$. A graph $G$ is monoholed if it is $\ell$-holed for some integer $\ell$. For any odd integer $\ell$, $\ell$-holed graphs are also even-hole free graphs (i.e. graphs with no holes of even length) and, for even integers $\ell$, it is known that $\ell$-holed graphs are perfect [27], that is, do not contain any odd holes or complements of an odd hole. Studies on even-hole free graphs [33] and perfect graphs [16] have yielded many powerful decomposition theorems and algorithmic techniques for fundamental problems such as Coloring, Independent Set, and Clique.
A complete structural description of $\ell$-holed graphs with $\ell\geq 7$ was obtained by Cook et al. [18]. Monoholed graphs can be recognised in polynomial time [18, 27]. The class of $5$-holed graphs was studied in the context of algebraic combinatorics and commutative algebra [35]. However, a complete description of $5$-holed graphs remains unknown [18]. Several subclasses of $5$-holed graphs have been studied in the literature [22, 30]. We show (without using the decomposition theorem from [18]) that monoholed graphs have bounded strong isometric path complexity.
Theorem 2.
The class of monoholed graphs is $U_{4}$-asymptotic minor-free. Hence, monoholed graphs have bounded strong isometric path complexity.
The above theorem implies that Isometric Path Cover admits a constant factor approximation algorithm on monoholed graphs. Some other algorithmic results on monoholed graphs are known. Maximum Independent Set can be solved in polynomial time on monoholed graphs [18]. Clique can be solved in polynomial time for $\ell$-holed graphs with $\ell\geq 5$, as there are only polynomially many maximal cliques in such graphs. The computational complexity of Coloring is open on these graphs. Already, optimally coloring the special case of rings is a difficult (but solved) problem [28].
Next, we consider two superclasses of monoholed graphs: perfect graphs and even-hole free graphs. Bipartite graphs (which form a subset of both these classes) have unbounded isometric path complexity. Indeed, it can be easily checked that the isometric path complexity of an $(n\times n)$-grid is $\Omega(n)$. Hence, perfect graphs have unbounded isometric path complexity. We show that the isometric path complexity of even-hole free graphs is also unbounded.
Theorem 3.
The class of even-hole free graphs with maximum degree at most 4 has unbounded isometric path complexity.
Graph operations: Quasi-isometry is a generalisation of biLipschitz maps that allows additive distortion. Embeddings with multiplicative (i.e. biLipschitz embeddings) or additive distortion are an area of active research [25, 24, 3, 5, 23, 21, 13]. Hyperbolic graphs are closed under quasi-isometry [26]. For a fixed graph $H$, $H$-asymptotic minor-free graphs are also closed under quasi-isometry [25]. Therefore, we investigate whether the class of graphs with bounded (strong) isometric path complexity is closed under quasi-isometry.
We prove that two popular quasi-isometric graph operations: graph powers and line graphspreserve strong isometric path complexity. For an integer $r\geq 1$ and a graph $G$, the $r^{th}$ power of $G$, denoted as $G^{r}$, is the graph with vertex set $V(G)$ where two vertices are adjacent in $G^{r}$ if they are at distance at most $r$ in $G$. The line graph of $G$, denoted as $L\left(G\right)$, has vertex set $E(G)$ and two vertices are adjacent in $L\left(G\right)$ if the corresponding edges share an end-vertex in $G$. Clearly, for a fixed integer $r$, any graph $G$ and its $r^{th}$ power ${G}^{r}$ are quasi-isometric. Similarly, any graph $G$ and its line graph $L\left(G\right)$ are also quasi-isometric. Line graphs and fixed powers of graphs are important in various areas of graph theory including metric graph theory [7]. We prove the following.
Theorem 4.
Let $G$ be a graph with isometric path complexity (resp. strong isometric path complexity) at most $k$. Then the following holds:
  1. (a)
    for every integer $r\geq 1$, the isometric path complexity (resp. strong isometric path complexity) of $G^{r}$ is at most $k(4r^{2}-2r)$;
  2. (b)
    the isometric path complexity (resp. strong isometric path complexity) of $L\left(G\right)$ is at most $23k+$ 2. Moreover, for every $t\geq 1$ there exists a graph $H_{t}$ with $ipco\left(H_{t}\right)=t$ and $ipco\left(L\left(H_{t}\right)\right)=2t-1$.
We do not known if the dependency on $r$ is needed in Theorem 4(a). Note that $r$-powers of chordal graphs (for even $r$) are also chordal, and chordal graphs have strong isometric path complexity at most 4 [11], so in this case, no dependency on $r$ is needed.
If two graphs $G$ and $H$ each contain cliques of equal size, the clique-sum of $G$ and $H$ is done from the disjoint union of $G$ and $H$ by identifying pairs of vertices in these two cliques to form a single shared clique. The resulting graph is $G\oplus H$. Clique-sums are important in graph minor theory. This operation is also studied in the context of decomposition theorems (e.g. for chordal graphs, universally-signable graphs [17] etc.). The clique-sum of two hyperbolic graphs (resp. $H$-asymptotic minor-free graphs) yields a hyperbolic graphs (resp. $H$-asymptotic minor-free graphs). However, it is known that the isometric path complexity is not closed under clique-sum [10]. We prove the contrary for strong isometric path complexity. For graphs $G_{1},G_{2},\ldots,G_{t}$, denote by $\bigoplus\limits_{i=1}^{t}G_{i}$ the graph obtained by taking pairwise clique sums of $G_{1},G_{2},\ldots,G_{t}$. (Note that it is not necessary that all graphs are identified using a single clique.)
Theorem 5.
Let $G_{1},G_{2},\ldots,G_{t}$ be graphs with $sipco\left(G_{i}\right)\leq k$ for all $i\in[t]$. Then the strong isometric path complexity of $G=\displaystyle\bigoplus\limits_{i=1}^{t}G_{i}$ is at most $183k+$ 18.
We suspect that the factor of $3$ is not needed in the bound of Theorem 5. Can it be improved to $k+c$, for some constant $c$? Note that there are graphs whose clique-sum has strictly larger strong isometric path complexity than the original graphs. For example, the strong isometric path complexity of chordal graphs (which can be constructed using clique-sums of complete graphs) is at most 4, whereas complete graphs have strong isometric path complexity 2.
Organization of the paper: In Section 2, we recall some definitions and some results. In Section 3 we prove Theorems 1, 2 and 3. In Section 4 we prove Theorems 4 and 5. We conclude in Section 5.
2 Preliminaries
In this section, we recall some definitions and some related observations. A sequence of distinct vertices forms a path $P$ if any two consecutive vertices are adjacent. Whenever we fix a path $P$ of $G$, we shall refer to the subgraph formed by the edges between the consecutive vertices of $P$. The length of a path $P$, denoted by $|P|$, is the number of its vertices minus one. A path is induced if there are no graph edges joining non-consecutive vertices. A path is isometric if it is a shortest path between its end-vertices. For two vertices $u,v$ of a graph $G$, $\mathsf{d}\left(u,v\right)$ denotes the length of an isometric path between $u$ and $v$. An isometric path $P$ is $v$-rooted if one of the end-vertices of $P$ is $v$.
For a path $P$ of a graph $G$ between two vertices $u$ and $v$, the vertices $V(P)\setminus\{u,v\}$ are internal vertices of $P$. A path between two vertices $u$ and $v$ is called a $(u,v)$-path. Similarly, we have the notions of isometric $(u,v)$-path and induced $(u,v)$-path. For a vertex $r$ of $G$ and a set $S$ of vertices of $G$, the distance of $S$ from $r$, denoted as $\mathsf{d}\left(r,S\right)$, is the minimum of the distance between any vertex of $S$ and $r$. For a subgraph $H$ of $G$, the distance of $H$ w.r.t. $r$ is $\mathsf{d}\left(r,V(H)\right)$. Formally, we have $\mathsf{d}\left(r,S\right)=\min\{\mathsf{d}\left(r,v\right)\colon v\in S\}$ and $\mathsf{d}\left(r,H\right)=\mathsf{d}\left(r,V(H)\right)$. For two sets $S$ and $T$, $\mathsf{d}\left(S,T\right)=\min\{\mathsf{d}\left(r,T\right)\colon r\in S\}$.
2.1 Asymptotic minors
For a positive integer $K$ and a graph $H$, a $K$-fat minor model of $H$ in a graph $G$ is a collection $\mathcal{M}=\left(B_{v}\colon v\in V(H)\right)\cup\left(P_{e}\colon e\in E(H)\right)$ of connected subgraphs of $G$ such that
  • $V(B_{v})\cap V(P_{e})\neq\emptyset$ whenever $v$ is an end-vertex of $e$ in $H$;
  • for any pair of distinct $X,Y\in\mathcal{M}$ not covered by the above condition, we have $\mathsf{d}_{G}\left(X,Y\right)\geq K$.
The subgraphs in $(B_{v}\colon v\in E(H))$ are the “branch sets” of the model. The subgraphs in $(P_{e}\colon e\in E(H))$ are the “branch paths” of the model. If $G$ contains a $K$-fat minor model of a graph $H$, then we say that $H$ is a $K$-fat minor of $G$.
Definition 2 ([25]).
A graph class $\cal G$ contains $H$ as an asymptotic minor if for every integer $K\geq 1$, there exists a graph $G\in\cal G$, such that $G$ contains $H$ as a $K$-fat minor.
We say that a graph class $\mathcal{G}$ is $H$-asymptotic minor-free if $\cal G$ does not contain $H$ as an asymptotic minor. In other words, $\cal G$ is $H$-asymptotic minor-free if there exists a constant $K$, depending only on $\mathcal{G}$, such that no graph in $\mathcal{G}$ contains a $K$-fat minor model of $H$.
Proposition 1.
For an integer $t\leq 1$, if a graph class $\cal G$ is $K_{2,t}$-asymptotic minor-free, then $\cal G$ is $U_{t}$-asymptotic minor-free.
Proof.
Assume for contradiction, that $\cal G$ is not $U_{t}$-asymptotic minor-free. This implies that for every integer $K$, there exists a graph $G\in\cal G$ containing a $K$-fat minor model $\mathcal{M}_{t}$ of $U_{t}$. Let $\mathcal{M}_{t}=\left(B_{w}\colon w\in V(U_{t})\right)\cup\left(P_{e}\colon e\in E(U_{t})\right)$. We shall construct a $\frac{K}{2}$-fat minor model of $K_{2,t}$. Let $u$ denote the universal vertex of $U_{t}$ and $B_{u}$ denote the corresponding branch set in $\mathcal{M}$. For each edge $uv\in E(U_{t})$, let $Q_{uv}\subset V(P_{uv})$ denote an induced path of length at least $K$ such that $Q_{uv}$ meets $B_{u}$ or $B_{v}$ only at its end-vertices. By definition, such a path always exists. For each edge $uv\in E(U_{t})$, let $x_{v}$ denote the vertex of $Q_{uv}$ which is at distance $\frac{K}{2}$ from both of end-vertices of $Q_{uv}$. Now we define the branch sets and branch paths of a $\frac{K}{2}$-fat minor model of a $K_{2,t}$. Define $B^{\prime}_{v}=\{x_{v}\}$. Also define $P^{1}_{v},P^{2}_{v}$ be the two paths obtained by deleting $x_{v}$ from $Q_{uv}$. Define $B^{\prime}_{0}=B_{u}$. Finally let $B^{\prime}_{1}$ be the graph induced by the vertices $\{V(B_{w})\colon w\neq u\in V(U_{t})\}\cup\{V(P_{e})\colon e=ww^{\prime},u\notin\{w,w^{\prime}\}\}$. It is easy to check that $\{B^{\prime}_{0},B^{\prime}_{1}\}\cup\{B^{\prime}_{v}\colon uv\in E(U_{t})\}\cup\{P^{1}_{v},P^{2}_{v}\colon uv\in E(U_{t})\}$ is a $\frac{K}{2}$-minor model of $K_{2,t}$. ∎
2.2 Isometric path complexity
For a graph $G$ and a vertex $r\in V(G)$, consider the following operations on $G$. First, remove all edges $xy$ from $G$ such that $\mathsf{d}\left(r,x\right)=\mathsf{d}\left(r,y\right)$. Let $G^{\prime}_{r}$ be the resulting graph. Then, for each edge $e=xy\in E(G^{\prime}_{r})$ with $\mathsf{d}\left(r,x\right)=\mathsf{d}\left(r,y\right)-1$, orient $e$ from $y$ to $x$. Let $\overrightarrow{G_{r}}$ be the directed acyclic graph formed after applying the above operation on $G^{\prime}$. Note that this digraph can easily be computed in linear time using a Breadth-First Search (BFS) traversal with starting vertex $r$.
The following definition in[10, 11] is inspired by the terminology of posets (as the graph $\overrightarrow{G_{r}}$ can be seen as the Hasse diagram of a poset). For a graph $G$ and a vertex $r\in V(G)$, two vertices $x,y\in V(G)$ are antichain vertices if there are no directed paths from $x$ to $y$ or from $y$ to $x$ in $\overrightarrow{G_{r}}$. A set $X$ of vertices of $G$ is an antichain set if any two vertices in $X$ are antichain vertices. A dual notion of isometric path complexity was defined in [10].
Let $r$ be a vertex of a graph $G$. For a subgraph $H$, $A_{r}\left(H\right)$ shall denote the maximum antichain set of $H$ in $\overrightarrow{G_{r}}$. Let $ipacc\left(\overrightarrow{G_{r}}\right)=\max\left\{|A_{r}\left(P\right)|\colon~{}P~{}\text{is an isometric path}\right\}.$ Define $ipacc\left(G\right)=\min\left\{ipacc\left(\overrightarrow{G_{r}}\right)\colon r\in V(G)\right\}.$
For a graph $G$ and a vertex $r$ of $G$, let $ipco\left(\overrightarrow{G_{r}}\right)$ denote the minimum integer $k$ such that any isometric path $P$ of $G$ can be covered by $k$ $r$-rooted isometric paths. Using Dilworth’s Theorem, one can prove the following important lemma and proposition.
Lemma 1 ([10]).
For any graph $G$ and vertex $r$, $ipco\left(\overrightarrow{G_{r}}\right)=ipacc\left(\overrightarrow{G_{r}}\right)$. Therefore, $ipco\left(G\right)=ipacc\left(G\right)$.
Thanks to Lemma 1, we may use one or the other notion depending on which is easier to use in a given proof. We shall use the following proposition. We provide its proof for completeness.
Proposition 2 ([10, 11]).
Let $G$ be a graph and $r$, an arbitrary vertex of $G$. Consider the directed acyclic graph $\overrightarrow{G_{r}}$, and let $P$ be an isometric path between two vertices $x$ and $y$ in $G$. Then $|P|\geq|\mathsf{d}\left(r,x\right)-\mathsf{d}\left(r,y\right)|+|A_{r}\left(P\right)|-1$.
Proof.
Orient the edges of $P$ from $y$ to $x$ in $G$. First, observe that $P$ must contain a set $E_{1}$ of oriented edges such that $|E_{1}|=|\mathsf{d}\left(r,y\right)-\mathsf{d}\left(r,x\right)|$ and for any $\overrightarrow{ab}\in E_{1}$, $\mathsf{d}\left(r,a\right)=\mathsf{d}\left(r,b\right)+1$. Let the vertices of the largest antichain set of $P$ in $\overrightarrow{G_{r}}$, i.e., $A_{r}\left(P\right)$, be ordered as $a_{1},a_{2},\ldots,a_{t}$ according to their occurrence while traversing $P$ from $y$ to $x$. For $i\in[2,t]$, let $P_{i}$ be the subpath of $P$ between $a_{i-1}$ and $a_{i}$. Observe that for any $i\in[2,t]$, since $a_{i}$ and $a_{i-1}$ are antichain vertices, there must exist an oriented edge $\overrightarrow{b_{i}c_{i}}\in E(P_{i})$ such that either $\mathsf{d}\left(r,b_{i}\right)=\mathsf{d}\left(r,c_{i}\right)$ or $\mathsf{d}\left(r,b_{i}\right)=\mathsf{d}\left(r,c_{i}\right)-1$. Let $E_{2}=\{b_{i}c_{i}\}_{i\in[2,t]}$. Observe that $E_{1}\cap E_{2}=\emptyset$ and therefore $|P|\geq|E_{1}|+|E_{2}|=|\mathsf{d}\left(r,y\right)-\mathsf{d}\left(r,x\right)|+|A_{r}\left(P\right)|-1$. ∎
3 Asymptotic minors and graphs with restricted holes
In Section 3.1 we show that $U_{t}$-asymptotic minor-free graphs have bounded strong isometric path complexity and prove Theorem 1. In Section 3.2 we show that monoholed graphs are $U_{4}$-asymptotic minor-free and prove Theorem 2. Finally, in Section 3.3 we show that even-hole free graphs have unbounded isometric path complexity and prove Theorem 3.
3.1 Asymptotic minors
In this section, we prove Theorem 1. Let $\mathcal{G}$ be a graph class that does not contain $U_{t}$ as an asymptotic minor, for some $t\geq 1$. Then, by the definition, there exists an integer $K$ and a graph $G\in\mathcal{G}$ such that $G$ does not contain a $K$-fat minor model of $U_{t}$. We can assume $K\geq 5$. The following lemma and Proposition 2 prove Theorem 1.
Lemma 2.
For any two integers $t\geq 2,K\geq 5$, if $G$ does not contain a $K$-fat minor model of $U_{t}$, then $ipacc\left(G\right)\leq(10K+9)t-1$.
We devote the remainder of this section to proving Lemma 2. For integers $t\geq 2,K\geq 5$, let $G$ be a graph that does not contain a $K$-fat minor model of $U_{t}$, such that $ipacc\left(G\right)\geq(10K+9)t$. First, we introduce some notation that will be used throughout the proof.
Let $r$ be any vertex of $G$. Note that $ipacc\left(\overrightarrow{G_{r}}\right)\geq(10K+9)t$. Hence, there exists an isometric path $P$ with $|A_{r}\left(P\right)|\geq(10K+9)t$. Assume that the two end-vertices of $P$ are $u$ and $v$, respectively. Notice that $r\notin\left\{u,v\right\}$. Let $A=A_{r}\left(P\right)$, $k=|A|$ and the elements of $A$ are ordered as $p_{1},p_{2},\ldots,p_{k}$ as they are encountered while traversing $P$ from $u$ to $v$. Let $a_{1}=p_{1}$, and for $i\in[2,t]$ let $n(i)=(10K+9)(i-1)$, $a_{i}=p_{n(i)}$. For convenience, we redefine $A=\left\{a_{i}\colon i\in[t]\right\}$. Recall that this set contains antichain vertices.
[…]
Figure 3: Illustration of the proof of Lemma 2. The shaded triangular shapes indicate the branch sets $Z_{1},Z_{2},\ldots,Z_{t}$, and subpaths of $P,Q_{1},Q_{2},\ldots,Q_{t}$ are used as branch paths to construct a $K$-fat minor model of $U_{t}$.
Input : A root vertex $r$, an isometric path $P$ whose end-vertices are distinct from $r$, a vertex $w\in V(P)$, and an end-vertex $u$ of $P$.
Output : An induced path.
1
2Let $Q$ be any $(r,w)$-isometric path in $G$;
3Let $Q^{\prime}$ be the maximal $(r,w^{\prime})$-subpath of $Q$ such that no internal vertex of $Q^{\prime}$ is adjacent to any vertex of $P$.
4if $w^{\prime}\in V(P)$ then
5       Return $Q^{\prime}$.
6else
7       Let $w^{\prime\prime}\in V(P)\cap N(w^{\prime})$ be the vertex such that no internal vertex of the $(u,w^{\prime\prime})$-subpath of $P$ is a neighbor of $w^{\prime}$.
8      Let $Q^{\prime\prime}$ be the path obtained by appending $w^{\prime\prime}$ to $Q^{\prime}$.
9      Return $Q^{\prime\prime}$.
10
Procedure 1 A procedure $\textsc{Connect}\left(r,P,w,u\right)$ to build pillars.
Finding special induced paths: Now we shall describe a general procedure to find special induced paths which will allow us to create branch paths and branch sets of a $K$-fat minor model of $U_{t}$ (and hence arrive to a contradiction). Consider Procedure 1. Let $Q_{i}$ denote the output of $\textsc{Connect}\left(r,P,a_{i},u\right)$ where $a_{i}\in A$. For each $i\in[t]$, observe that one end-vertex of $Q_{i}$ lies on $P$. Let us denote this end-vertex as $b_{i}$ and its neighbor in $Q_{i}$ as $b^{\prime}_{i}$. We prove the following claims that establish the special properties of these paths.
Claim 5.1.
For each $i\in[t]$, the following holds:
  1. (a)
    the path $Q_{i}$ is an induced path;
  2. (b)
    no vertex of $V(Q_{i})\setminus\left\{b_{i},b^{\prime}_{i}\right\}$ is adjacent to a vertex of $P$;
  3. (c)
    the $(r,b^{\prime}_{i})$-subpath of $Q_{i}$ is a subpath of an $(r,a_{i})$-isometric path.
Proof.
To prove (a) fix any $i\in[t]$. The path $Q_{i}$ is the path returned by $\textsc{Connect}\left(r,P,a_{i},u\right)$. If $\textsc{Connect}\left(r,P,a_{i},u\right)$ returned from Step-$4$ of Procedure 1, then $Q_{i}$ is an $(r,w_{i})$-isometric path and therefore is an induced path. Otherwise, $\textsc{Connect}\left(r,P,a_{i},u\right)$ must have returned from Step-$8$ of Procedure 1. In this case, $Q^{\prime}_{i}$ is constructed by taking a subpath of a $(r,a_{i})$-isometric path and appending to it one more vertex from the path $P$. By construction (see Step-$1$ and Step-$6$), $Q^{\prime}_{i}$ remains an induced path. Both $(b)$ and $(c)$ follow from the construction. ∎
The properties proved in the above claim will allow us to use the paths $\{Q_{i}\colon i\in[t]\}$ to construct branch paths and branch sets of $U_{t}$, and reach a contradiction. Note that the end-vertex of $Q_{i}$ that lies on $V(P)$ (i.e. $b_{i}$) may not be the same as $a_{i}$ which was initially used (in $\textsc{Connect}\left(r,P,a_{i},u\right)$) to construct $Q_{i}$. We prove the following claim.
Claim 5.2.
For each $i\in[t]$, let $X_{i}$ denote the $(a_{i},b_{i})$-subpath of $P$. Then $|A_{r}\left(X_{i}\right)|\leq 3$.
Proof.
Assume for contradiction that there exists an integer $i\in[t]$ such that $|A_{r}\left(X_{i}\right)|\geq 4$. Hence due to Proposition 2, we have that $\mathsf{d}\left(b_{i},a_{i}\right)\geq 3+|\mathsf{d}\left(r,b_{i}\right)-\mathsf{d}\left(r,a_{i}\right)|$. Let $R$ denote the $(r,b^{\prime}_{i})$-subpath of $Q_{i}$. Due to 5.1(c) we know that there is an $(r,a_{i})$-isometric path $S$ such that $V(R)\subseteq V(S)$. Hence $b^{\prime}_{i}\in V(S)$ and let $S^{\prime}$ denote the $(b^{\prime}_{i},a_{i})$-subpath of $S$. Observe that appending $b_{i}$ to $S^{\prime}$ gives a path from $b_{i}$ to $a_{i}$. Hence, $\mathsf{d}\left(b_{i},a_{i}\right)\leq 2+|\mathsf{d}\left(r,b_{i}\right)-\mathsf{d}\left(r,a_{i}\right)|$. But this is a contradiction. ∎
Claim 5.3.
For $\{i,j\}\subseteq[t]$, let $R_{ij}$ denote the $(b_{i},b_{j})$-subpath of $P$. Then $|A_{r}\left(R_{ij}\right)|\geq(10K+3)$.
Proof.
Without loss of generality assume $i<j$. Let $S$ denote the $(a_{i},a_{j})$-subpath of $P$. Observe that $|A_{r}\left(S\right)|\geq 10K+9$. Let $X_{i}$ and $X_{j}$ denote the $(a_{i},b_{i})$-subpath of $P$. Due to 5.2, we have that $|A_{r}\left(X_{i}\right)|+|A_{r}\left(X_{j}\right)|\leq 6$. Now consider the case when both $b_{i}$ and $b_{j}$ lies on the path $S$. In this case, $S$ is obtained by concatenating $X_{i}$, $R_{ij}$ and $X_{j}$. Hence, we have the following.
$|A_{r}\left(X_{i}\right)|+|A_{r}\left(R_{ij}\right)|+|A_{r}\left(X_{j}\right)|$ $\geq|A_{r}\left(S\right)|$
$|A_{r}\left(R_{ij}\right)|+6$ $\geq 10K+9$
$|A_{r}\left(R_{ij}\right)|$ $\geq 10K+3$
In the remaining cases i.e. when $b_{i}$ or $b_{j}$ does not lie on the path $S$, similar arguments prove the claim. ∎
The above claim implies that the $(b_{i},b_{j})$-subpath of $P$, with $i\neq j$, necessarily contains many anti-chain vertices.
Claim 5.4.
For an integer $i\in[t]$, let $c\in V(P)$ be a vertex such that the $(c,b_{i})$-subpath of $P$ contain at least $32K+$ 3 antichain vertices. Then for any vertex $d\in V(Q_{i})$, $\mathsf{d}\left(c,d\right)\geq K$.
Proof.
Suppose $\mathsf{d}\left(c,d\right)<K$. Then $\mathsf{d}\left(c,b_{i}\right)\leq 2k+1+|\mathsf{d}\left(r,b_{i}\right)-\mathsf{d}\left(r,c\right)|$. But this contradicts Proposition 2 which implies that $\mathsf{d}\left(c,b_{i}\right)\geq 2K+2+|\mathsf{d}\left(r,b_{i}\right)-\mathsf{d}\left(r,c\right)|$. ∎
For each $i\in[t]$, let $Q^{\prime}_{i}$ denote the maximal $(b_{i},c_{i})$-subpath of $Q_{i}$ such that $\mathsf{d}\left(Q^{\prime}_{i},Q_{j}\right)\geq K$ for some $i\neq j,j\in[t]$. Now we prove a lower bound on the length of $Q^{\prime}_{i}$, for each $i\in[t]$. We prove the following final claim before providing the explicit constructions of the branch paths and branch sets.
Claim 5.5.
For each $i\in[t]$, the path $Q^{\prime}_{i}$ has length at least $18K+$ 1.
Proof.
Assume for contradiction that there exists an integer $i\in[t]$ such that the length of $Q^{\prime}_{i}$ is at most $8K$. Let $c^{\prime}_{i}$ be the vertex adjacent to $c_{i}$ in $Q_{i}$ which is not a vertex of $Q^{\prime}_{i}$. Let $j\neq i$ be the minimum index such that $c^{\prime}_{i}$ lies within a distance of at most $2K$ from a vertex $d_{j}\in V(Q_{j})$. Let $T_{ij}$ denote a $(c^{\prime}_{i},d_{j})$-isometric path in $G$. Now, observe that the path $X=Q^{\prime}_{i}\oplus\{c^{\prime}_{i}\}\oplus T_{ij}\oplus Q^{\prime}_{j}$ is a $(b_{i},b_{j})$-path. Since $\mathsf{d}\left(b_{i},b_{j}\right)$ is at most the length of $X$ we have that
$\mathsf{d}\left(b_{i},b_{j}\right)\leq 10K+|\mathsf{d}\left(r,b_{j}\right)-\mathsf{d}\left(r,b_{i}\right)|$
Let $R_{ij}$ denotes the $(b_{i},b_{j})$-subpath of $P$. From 5.3 it follows that $|A_{r}\left(R_{ij}\right)|\geq 10K+3$. Now from Proposition 2 it follows that
$\mathsf{d}\left(b_{i},b_{j}\right)\geq 10K+2+|\mathsf{d}\left(r,b_{j}\right)-\mathsf{d}\left(r,b_{i}\right)|$
which is a contradiction. ∎
Constructing branch paths: Let $\mathcal{Q}=\{Q^{\prime}_{i}\colon i\in[t]\}$. Let $X$ be the $(b_{i},b_{i+1})$-subpath of $P$ and let $D_{i}$ denote the maximum subpath of $X$ such that both components of the graph $E(X)\setminus E(D_{i})$ contain $24K+$ 2 antichain vertices. Observe that $D_{i}$ always exists (due to 5.3). Let $\mathcal{D}=\{D_{i}\colon i\in[t-1]\}\}$. Moreover, each $D_{i}$ contains at least $K+3$ antichain vertices. We shall use this fact later.
Claim 5.6.
For each $i\in[t-1]$, the length of $D_{i}$ is at least $K+3$.
We now show that the paths in $\mathcal{Q}\cup\mathcal{D}$ can be used as branch paths.
Claim 5.7.
For any two paths $X,Y\in\mathcal{Q}\cup\mathcal{D}$, $\mathsf{d}\left(X,Y\right)\geq K$.
Proof.
If $X,Y\in\mathcal{C}$, then we are done by the definition of $\mathcal{Q}$. Suppose $X$ is $Q^{\prime}_{i}$ for some $i\in[t]$ and $Y=D_{j}$ for some $j\in[t-1]$. Assume that $\mathsf{d}\left(X,Y\right)\leq K$ and $y\in Y$ be a vertex such that $\mathsf{d}\left(y,X\right)<K$. Note that $y\in V(P)$ and the $(b_{i},y)$-subpath of $P$, say $P^{\prime}$, contains at least $32K+$ 3 antichain vertices. Hence, we have a contradiction due to 5.4. ∎
Constructing branch sets: For each $i\in[t]$, let $c^{\prime}_{i}$ denote the end-vertex of $Q^{\prime}_{i}$ that is closest to $r$, and let $T_{i}$ denote the $(r,c^{\prime}_{i})$-subpath of $Q_{i}$. Let $T$ denote the set $\bigcup\limits_{i\in I}V(T_{i})$. Clearly, the subgraph induced by $T$ is connected and $T$ will be the branch set corresponding to the universal vertex of $U_{t}$. Below, we construct the branch sets corresponding to the other vertices of $U_{t}$. We shall use the paths in $\mathcal{D}$ (defined above) to construct these branch sets.
For each $i\in[t]$, let $X_{i}$ denote the $(a_{i},b_{i})$-subpath of $P$ and $P_{i}=V(X_{i})\cup\{b^{\prime}_{i}\}$. Observe that the subgraph induced by $P_{i}$ is connected. For $i\in[t-1]$, let $R_{i}$ denote the $(b_{i},b_{i+1})$-subpath of $P$. Let $Y_{1}$ be the vertices of the connected component of $R_{1}\setminus D_{1}$ that contains $b_{1}$. For $i\in[2,t-1]\}$, let $Y_{i}$ be the union of the vertices of the connected components of $E(R_{i-1})\setminus E(D_{i-1})$ and $E(R_{i+1})\setminus E(D_{i+1})$ that contain $b_{i}$. Finally, let $Y_{t}$ be the vertices of the connected component of $R_{t-1}\setminus D_{t-1}$ that contains $b_{t}$. For $i\in[t]$, let $Z_{i}=Y_{i}\cup P_{i}$. Note that the subgraph induced by $Z_{i}$ is connected and $Z_{i}\subseteq V(Q^{\prime}_{i})\cup V(P)$. We have the following claims.
Claim 5.8.
For any $i\in[t]$, $\mathsf{d}\left(T,Z_{i}\right)\geq K$.
Proof.
Suppose that there exists an integer $i\in[t]$, such that $\mathsf{d}\left(x,y\right)<K$ with $x\in T$ and $y\in Z_{i}$. The definition of $T$ implies that $x\in Q_{j}$ for some $j\in[t]$. 5.5 immediately implies that $y\in V(P)\cap Z_{j}$. Now if $i\neq j$, the $(y,b_{j})$-subpath of $P$ contains at least $34K+$ 3 antichain vertices, which is a contradiction due to 5.4. Now consider the case when $i=j$. Let $c^{\prime}_{i}$ denote the end-vertex of $Q^{\prime}_{i}$ that is different from $b_{i}$. From the definition of $T$ (and $T_{i}$), it follows that $\mathsf{d}\left(r,x\right)\leq\mathsf{d}\left(r,c^{\prime}_{i}\right)$.
Let $d^{\prime}_{i}$ denote the vertex of $Q_{i}$ which is adjacent to $c^{\prime}_{i}$ but does not lie in $Q^{\prime}_{i}$. The definition of $Q^{\prime}_{i}$ implies that there exists an integer $k\in[t]\setminus\{i\}$, such that $\mathsf{d}\left(c^{\prime}_{i},Q^{\prime}_{k}\right)<K$. Let $d^{\prime}_{k}$ denote a vertex of $Q^{\prime}_{k}$ with $\mathsf{d}\left(d^{\prime}_{i},d^{\prime}_{k}\right)<K$. Observe the following:
$\mathsf{d}\left(b_{k},y\right)$ $\leq 1+\mathsf{d}\left(b^{\prime}_{k},d^{\prime}_{k}\right)+\mathsf{d}\left(d^{\prime}_{k},d^{\prime}_{i}\right)+\mathsf{d}\left(d^{\prime}_{i},x\right)+\mathsf{d}\left(x,y\right)$
$\leq 1+|\mathsf{d}\left(r,b^{\prime}_{k}\right)-\mathsf{d}\left(r,d^{\prime}_{k}\right)|+K+|\mathsf{d}\left(r,d^{\prime}_{i}\right)-\mathsf{d}\left(r,x\right)|+K$
$\leq 4K+1+|\mathsf{d}\left(r,b^{\prime}_{k}\right)-\mathsf{d}\left(r,y\right)|$
$\leq 4K+|\mathsf{d}\left(r,b_{k}\right)-\mathsf{d}\left(r,y\right)|$
Now observe that the $(b_{k},y)$-subpath of $P$ has at least $24K+$ 2 antichain vertices. Therefore, Proposition 2 implies that $\mathsf{d}\left(y,b_{k}\right)\geq 4K+1+|\mathsf{d}\left(r,b_{k}\right)-\mathsf{d}\left(r,y\right)|$. But this is is a contradiction. ∎
Claim 5.9.
For any $i,j\in[t]$ with $i<j$, $\mathsf{d}\left(Z_{i},Z_{j}\right)\geq K$.
Proof.
This follows from 5.6 and the definition of $Z_{i}$. ∎
Now we show that the branch sets constructed above are sufficiently far away from the branch paths as required.
Claim 5.10.
For any $i\in[t-1]$, $\mathsf{d}\left(D_{i},T\right)\geq K$.
Proof.
Suppose there is a vertex $x\in D_{i}$ and $y\in T$ such that $\mathsf{d}\left(x,y\right)<K$. The definition of $T$ implies that $y\in V(Q_{j})$ for some $j\in[t]$. Since the $(b_{j},x)$-subpath of $P$ contains at least $34K+$ 3 antichain vertices, we have a contradiction due to 5.4. ∎
Using similar arguments as above, we can prove the following.
Claim 5.11.
For any $i\in[t-1],j\in[t]$ with $(j-i)\notin\{0,1\}$, we have $\mathsf{d}\left(D_{i},Z_{j}\right)\geq K$.
Completion of the proof of the Lemma 2: Now consider the sets $T,Z_{1},\ldots,Z_{t}$, $\mathcal{Q}=\{Q^{\prime}_{1},Q^{\prime}_{2},\ldots,Q^{\prime}_{t}\}$, and $\mathcal{D}=\{D_{1},D_{2},\ldots,D_{t-1}\}$. Observe that
  • for each $i\in[t]$, $T\cap Q^{\prime}_{i}\neq\emptyset$;
  • $D_{1}\cap Z_{1}\neq\emptyset,D_{t-1}\cap Z_{t}\neq\emptyset$; and
  • for $i\in[2,t-1],j\in[t-1],j-i\in\{0,1\},$ $Z_{i}\cap D_{j}\neq\emptyset$.
Now 5.7, 5.8, 5.9, 5.10, and 5.11 imply that the sets $T,Z_{1},\ldots,Z_{t}$ and $\mathcal{Q}\cup\mathcal{D}$ form a $K$-fat minor model of $U_{t}$, a contradiction.
3.2 Monoholed graphs
Now, we shall prove Theorem 2. We need the definitions of Truemper configurations [34]. A theta is a graph made of three vertex-disjoint induced paths $P_{1}=a\ldots b$, $P_{2}=a\ldots b$, $P_{3}=a\ldots b$ of lengths at least 2, and such that no edges exist between the paths except the three edges incident to $a$ and the three edges incident to $b$. A pyramid is a graph made of three induced paths $P_{1}=a\ldots b_{1}$, $P_{2}=a\ldots b_{2}$, $P_{3}=a\ldots b_{3}$, two of which have length at least $2$, that are vertex-disjoint (except at $a$), such that $b_{1}b_{2}b_{3}$ is a triangle, and no edges exist between the paths except those of the triangle and the three edges incident to $a$. A prism is a graph made of three vertex-disjoint induced paths $P_{1}=a_{1}\ldots b_{1}$, $P_{2}=a_{2}\ldots b_{2}$, $P_{3}=a_{3}\ldots b_{3}$ of lengths at least $1$, such that $a_{1}a_{2}a_{3}$ and $b_{1}b_{2}b_{3}$ are triangles and no edges exist between the paths except those of the two triangles.
A $3$-path configuration is a theta, prism or a pyramid. A $3$-path configuration is balanced if the lengths of all the paths in the respective definitions are the same; otherwise, it is unbalanced. The length of a balanced 3-path configuration is the length of the paths in the respective definitions. Notice that all $3$-path configurations in a monoholed graph must be balanced and of the same length. Now, we will be done by proving the following lemma.
Lemma 3.
Let $G$ be a graph such that all $3$-path configurations contained in $G$ are balanced and of the same length. Then $G$ does not contain a $10$-fat minor model of $U_{4}$.
Proof.
[…]
(a) (b)
Figure 4: (a)The graph $U_{4}$. (b) Illustration of the proof of Theorem 2. The thick, black, closed curve indicates the induced cycle $D$ (in 5.12).
Assume for contradiction that $G$ contains a $K$-fat minor model of $U_{4}$ with $K\geq 10$. Let the vertices of $U_{4}$ be denoted by $a_{0},a_{1},a_{2},a_{3},a_{4}$ as shown in Figure 4(a). Unless stated explicitly, all addition operations in this proof will be taken modulo $5$. Let $B_{i}$ denote the branch set corresponding to the vertex $a_{i}$ and $C_{i}^{j}$ denote the branch path corresponding to the edge between $a_{i}$ and $a_{j}$. See Figure 4(b) for illustrations. For the sake of readability, we shall use the same notations to indicate the subgraphs induced by the branch sets and branch paths. Let $Z=\{(0,1),(0,2),(0,3),(0,4),(1,2),(2,3),(3,4)\}$. A witness of a branch path $C_{i}^{j},(i,j)\in Z$ is an induced path $P$ such that $V(P)\subseteq C_{i}^{j}$, one end-vertex of $P$ lies in $B_{i}$, and the other end-vertex of $P$ lies in $B_{j}$. For $0\leq i\leq 4$, let $W_{i}$ be a witness of $C_{i}^{i+1}$ of shortest possible length, and let $u_{i}\in B_{i},v_{i+1}\in B_{i+1}$ be the end-vertices of $W_{i}$. For $0\leq i\leq 4$, $X_{i}$ denote a path of shortest possible length between $u_{i}$ and $v_{i}$ such that $V(X_{i})\subseteq B_{i}$. Let $H$ be the subgraph of $G$ induced by $\bigcup\limits_{i=0}^{4}V(X_{i})\cup V(W_{i})$.
Claim 5.12.
The graph $H$ contains an induced cycle $D$ such that for each $0\leq i\leq 4$, both $V(B_{i})\cap V(D)$ and $V(C_{i}^{i+1})\cap V(D)$ induce paths of lengths at least $K-2$.
Proof.
The minimality of $W_{i}$ implies that only the end-vertices of $W_{i}$ or their neighbors (in $W_{i}$) can be adjacent to some vertices of $X_{i}$ or $X_{i+1}$. Hence, there must exist a subpath of $W_{i}$, say $W_{i}^{\prime}$, of length at least $K-2$ such that no internal vertex of $W_{i}^{\prime}$ is adjacent to a vertex of $X_{i}$ or $X_{i+1}$. Since the distance between two witnesses (corresponding to two branch paths) is at least $K$, observe that there must exist a subpath of $X_{i}$, say $X^{\prime}_{i}$, of length at least $k-2$ none of whose internal vertex is adjacent to the vertices of $W_{i}$ or $W_{i+1}$. Now, using $W_{i}^{\prime}$ and $X^{\prime}_{i}$, $0\leq i\leq 4$, it is possible to construct an induced cycle $D$ satisfying the claim. ∎
Let $D$ be the induced cycle whose existence was proved in 5.12. For $i\in\{0,1,2,3,4\}$, let $D_{i}=B_{i}\cap D$. Let $Q$ be an induced path of the shortest possible length between a vertex of $D_{2}$ and $D_{0}$ such that $V(Q)\subseteq B_{0}\cup C_{0}^{2}\cup B_{2}$. (Since $B_{0},C_{0}^{2},B_{2}$ are connected induced subgraphs, $Q$ always exists). Observe that $Q$ and $D$ together creates a balanced $3$-path configuration $P_{1}$, $P_{2}$, $P_{3}$ such that $V(P_{1})\cap V(W_{0})\neq\emptyset,V(P_{2})\cap C_{0}^{2}\neq\emptyset$ and $V(P_{3})\cap V(W_{4})\neq\emptyset$. For $i\in[3]$, let $z_{i}$ (resp. $z^{\prime}_{i})$ denote the end-vertex of $P_{i}$ which is adjacent to some vertices of $D_{0}$ (resp. $D_{2})$. Let $K^{\prime}$ be the lengths of the three paths. Observe that the vertices of the paths $P_{2}$ and $P_{1}$ (resp. $P_{3}$) induces a cycle $F$ (resp. $F^{\prime}$) of order at most $22K^{\prime}+$ 2 which also contains $D_{1}$ (resp. $D_{3}$).
Now consider an induced path $R$ of the shortest possible length between a vertex of $D_{3}$ and $D_{0}$ such that $V(R)\subseteq B_{0}\cup C_{0}^{3}\cup B_{3}$. Let $x$ and $y$ denote the end-vertices of $R$ that lie on $D_{3}$ and $D_{0}$, respectively. Let $x^{\prime},y^{\prime}$ be the neighbours of $x$ and $y$, respectively and $R^{\prime}$ be the $(x^{\prime},y^{\prime})$-subpath of $R$. The minimality of $R$ implies that no internal vertex of $R^{\prime}$ is adjacent to any vertex of $D_{0}$ and $D_{3}$. Furthermore, we prove the following claim.
Claim 5.13.
No vertex of $R^{\prime}$ is adjacent to any vertex of $F^{\prime}$.
Proof.
Suppose $w$ is the vertex of $R^{\prime}$, which is closest to $x^{\prime}$ and is also adjacent to some vertex $w^{\prime}$ of $F^{\prime}$. If $w^{\prime}\in P_{3}$, then $F^{\prime}$ together with the $(w,x^{\prime})$-subpath must create a $3$-path configuration which is either unbalanced or has length strictly less than $K^{\prime\prime}$, a contradiction. For similar reasons, $w^{\prime}$ cannot be $z_{2}$. Now consider the case when $w^{\prime}$ is an internal vertex of $P_{2}$. Let $U$ be the maximal subpath of $P_{3}$ whose one end-vertex is $z^{\prime}_{3}$, and no internal vertex of $U$ is a neighbor of $x^{\prime}$. Observe that $U$ together with the $(w,x^{\prime})$-subpath of $R^{\prime}$ creates an induced path $U^{\prime}$ whose (only) end-vertices are adjacent to the vertices of $F$ (the induced cycle created by $P_{2}$ and $P_{1}$). Since $w^{\prime}$ is an internal vertex, $F$ together with $U^{\prime}$ must create a $3$-path configuration which is either unbalanced or has a length strictly less than $K^{\prime}$, a contradiction. ∎
Let $S$ be the $(y,z_{3})$-subpath of $D_{0}$ and let $S^{\prime}$ be the subpath of $S$ of minimum possible length such that one end-vertex of $S^{\prime}$ is adjacent to $y^{\prime}$ and the other end-vertex is adjacent to $z_{2}$. Let $R^{\prime\prime}$ denote the path induced by $V(R^{\prime})\cup V(S^{\prime})$. The induced cycle $F^{\prime}$ along with $R^{\prime\prime}$ creates an unbalanced $3$-path configuration, a contradiction. ∎
3.3 Even-hole free graphs
$r_{1}$$a_{1}$$c_{1}$$b_{1}$$r_{2}$$a_{2}$$c_{2}$$b_{2}$$r_{3}$$a_{3}$$c_{3}$$b_{3}$$r_{4}$$a_{4}$$c_{4}$$b_{4}$
(a) (b)
Figure 5: Proof of Theorem 3. (a) The graph $B_{4}$. (b) The graph $G_{4}$.
It is enough to prove Theorem 3 for each integer $k\geq 4$. For an integer $k\geq 4$, let $B_{k}$ be the graph obtained by taking a triangle $T$ and attaching to one of its vertices a path of length $k$. Let $F_{1}$ denote a copy of $B_{k}$. Let the three vertices of the only triangle of $F_{1}$ be denoted as $a_{1},b_{1},c_{1}$ where the degree of $a_{1}$ is 3, and the only vertex of degree 1 is denoted as $r_{1}$. Let $G_{1}=F_{1}$. For $i=2,\ldots,k$ we construct $G_{i}$ as follows.
  • Let $F_{i}$ denote a copy of $B_{k}$. Let the three vertices of the only triangle of $F_{i}$ are denoted as $a_{i},b_{i},c_{i}$ where the degree of $a_{i}$ is three and the only vertex of degree one is denoted as $r_{i}$.
  • Connect $r_{i}$ and $r_{i-1}$ using a path $P_{i}$ of length $k$. Connect $c_{i}$ and $b_{i-1}$ using a path $Q_{i}$ of length $k+1$. The resulting graph is $G_{i}$.
Let $R=\{r_{i}\colon i\in[k]\}$. We prove the following lemma.
Lemma 4.
The graph $G_{k}$ does not contain any even hole.
Proof.
Let $C$ be an induced cycle of length at least $4$. It is clear from the construction that $C$ will contain at least one vertex from $R$. Moreover, if $r_{i}\in V(C),r_{t}\in V(C)$, then for any $i<j<t$, $r_{j}\in V(C)$. Let $i$ (resp. $t$) be the minimum (resp. maximum) index such that $r_{i}\in V(C)$ (resp. $r_{t}\in V(C)$). Hence, for each $i<j\leq t$, the path $P_{j}$ is also part of $C$. It is also easy to verify that for each $i<j\leq t$, the path $Q_{j}$ is also part of $C$. Let $X_{i}$ (resp. $X_{t}$) be the path between $r_{i}$ and $b_{i}$ (resp. $r_{t}$ and $c_{t}$). Observe that
$V(C)=V(X)\cup V(X^{\prime})\displaystyle\bigcup\limits_{i<j\leq t}\left(V(P_{j})\cup V(Q_{j})\right)\displaystyle\bigcup\limits_{i<j<t}\{b_{j},c_{j}\}$
Hence the length of $C$ is $(t-i)k+(t-i)(k+1)+2(k+1)+(t-i)-1$ which is always odd. ∎
Lemma 5.
Let $r=r_{k},G=G_{k}$. Then $ipco\left(\overrightarrow{G_{r}}\right)\geq k$.
Proof.
First, observe that for each $i\in[k]$, $\mathsf{d}\left(r,b_{i}\right)=\mathsf{d}\left(r,c_{i}\right)$. Therefore, if there is an isometric path in $G$ that contains both $b_{i}$ and $c_{i}$, for some $i\in[k]$, then two distinct $r$-rooted isometric paths are needed to cover $b_{i}$ and $c_{i}$. Moreover, it is clear from the construction that for $i<j\in[k]$, there is no directed path from $b_{j}$ to $b_{i}$ in $\overrightarrow{G_{r}}$. Now consider an isometric path $X$ in $G$, and let $t$ be the number of edges of the form $b_{i}c_{i}\in E(X)$. In other words, $t=|\{i\in[k]\colon\{b_{i},c_{i}\}\subseteq V(X)\}|$. Then the above arguments imply that $|A_{r}\left(X\right)|\geq t$. Now we will be done by showing that there is an isometric path that contains at least $k$ many edges of the form $b_{i}c_{i}$.
Let $Q$ be the induced path between $b_{k}$ and $c_{1}$ induced by the vertices of $\left(\displaystyle\bigcup\limits_{i=2}^{k}V(Q_{i})\right)\cup\left(\displaystyle\bigcup\limits_{i=1}^{k}\{b_{j},c_{j}\}\right)$. The length of $Q$ is $k^{2}+k-1$. We claim that $Q$ is an isometric path between $c_{1}$ and $b_{k}$. To prove this, consider a path $Q^{\prime}$ between $c_{1}$ and $b_{k}$ that is different from $Q$. Notice that $Q^{\prime}$ must contain at least one vertex from $R$ and $i$ be the minimum index such that $r_{i}\in R$. Observe that $\mathsf{d}\left(b_{k},r_{i}\right)\geq(k-i+1)k+1$ and $\mathsf{d}\left(r_{i},c_{1}\right)\geq(i+1)k+1$. Therefore, the length of $Q^{\prime}$ is at least $(k+2)k+2$, which is larger than $Q$. Therefore, $Q$ is isometric and contains all edges in $\{b_{i}c_{i}\colon i\in[k]\}$. ∎
Completion of the proof of Theorem 3: Consider two disjoint copies of the graph $G_{k}$, and identify the vertex $r_{k}$ from each copy. Let $Z_{k}$ be the resulting graph. Let $G=G_{k}$ and $r=r_{k}$. Observe that $Z_{k}$ is even-hole free, due to Lemma 4 and the fact that identifying one vertex from two disjoint even-hole free graph results in an even-hole free graph. Moreover, $ipco\left(Z_{k}\right)\geq ipco\left(\overrightarrow{G_{r}}\right)\geq k$, due to Lemma 5.
4 Graph operations
4.1 Graph powers
Let $G$ be a graph with $ipco\left(G\right)\leq k$ and $r\geq 2$ be a fixed integer. Let $H=G^{r}$. We will use the following simple observation.
Observation 1.
Let $P$ be a red $(u,v)$-isometric path of length $\ell\geq 1$ in $H$. Then the distance between $u$ and $v$ in $G$ is at least $r(\ell-1)+1$.
We prove the following lemma.
Lemma 6.
The vertices of any $v$-rooted isometric path $Q$ of $G$ can be covered by $r$ many $v$-rooted isometric paths in $H$.
Proof.
Let the vertices of $Q$ be ordered as $v_{1}=v,v_{2},\ldots,v_{\ell}$ as they are encountered while traversing $Q$ from $v$. Observe that for $i\in[r]$, the vertices in the set $P_{i}=\{v_{j}\colon j=i+1+xr,x\in\mathbb{N},xr\leq\ell\}$ induce an isometric path in $H$. Since $v$ and $v_{i}$ with $i\in[2,r]$ are adjacent in $H$, we have that $Q_{i}=P_{i}\cup\{v\}$ is also an isometric path in $H$. Now, the set $\mathcal{Q}=\{Q_{i}\colon i\in[r]\}$ covers all the vertices of $Q$. ∎
Now color the edges of $H$ as follows: an edge $e\in E(H)$ is black if $e\in E(G)$; otherwise $e$ is red. An isometric path $P$ of $H$ is red if all edges of $P$ are red.
Input : A graph $G$ and a path $P$ of $G$.
Output : Vertex-minimal non-isometric subpaths of $P$.
1
2if $P$ is isometric in $G$ then
3      Return $\emptyset$.
4else
5       Let $Q$ be any vertex-minimal non-isometric subpath of $P$;
6      $P_{1}$ and $P_{2}$ be the two paths in the subgraph $E(P)\setminus E(Q)$;
7      $Q_{1}=\textsc{Find-Path}(G,P_{1})$
8      $Q_{2}=\textsc{Find-Path}(G,P_{2})$
9      Return $\{Q\}\cup Q_{1}\cup Q_{2}$;
10
Procedure 2 A procedure Find-Path($G,P$) to find vertex-minimal non-isometric paths.
Lemma 7.
Any isometric path of $H$ contains at most one black edge.
Proof.
Assume for contradiction that there is a path $P$ in $H$ that has at least two black edges $u_{1}v_{1}$ and $u_{2}v_{2}$. Let the vertices $u_{1},v_{1},u_{2},v_{2}$ appear in that order while traversing $P$ from one end-vertex to the other. Also, assume that the subpath $P^{\prime}$ of $P$ between $u_{1}$ and $v_{2}$ does not contain any other black edge. Clearly, $u_{2}\neq v_{1}$. (Otherwise, the distance between $u_{1}$ and $v_{2}$ is at most 2 in $G$, which implies that $u_{1},v_{2}$ would be adjacent in $H$, contradicting that $P$ is isometric.) Hence, $P^{\prime}$ has at least one edge, and $P^{\prime}$ is a red isometric path in $H$ with $v_{1}$ and $u_{2}$ as its end-vertices. Let $x_{1}=v_{1},x_{2},\ldots,x_{\ell}=u_{2}$ be the vertices of $P^{\prime}$ as they are encountered while traversing $P^{\prime}$ from $v_{1}$ to $u_{2}$. For each $i\in[\ell-1]$, let $P_{i}$ denote an $(x_{i},x_{i+1})$-isometric path of length at most $k$ in $G$. (Since $x_{i}x_{i+1}$ is a red edge, such a path always exists.) For each $i\in[\ell-1]$, let $y_{i}$ denote the vertex of $P_{i}$ that is adjacent to $x_{i}$. Observe that, for each $i\in[\ell-2]$, $y_{i}$ is adjacent to $y_{i+1}$ in $H$. Moreover, $u_{1}$ is adjacent to $y_{1}$ in $H$ and $y_{\ell-1}$ is adjacent to $v_{2}$ in $H$. Hence, $u_{1},y_{1},y_{2},\ldots,y_{\ell-1},v_{2}$ induces a path of length $\ell$ in $H$. But this contradicts the fact that $u_{1},v_{1}=x_{1},\ldots,x_{\ell}=v_{2},u_{2}$ induces an isometric subpath of $P$ with length $\ell+1$. ∎
Lemma 8.
The vertices of any red isometric path of $H$ can be covered with $12r-$ 1 many isometric paths of $G$.
Proof.
Let $P$ be a red isometric path of $H$ with $|E(P)|=\ell$. Let $u_{1},u_{2},\ldots,u_{\ell+1}$ be the vertices of $P$ ordered as they are traversed from one end-vertex of $P$ to the other. Observe that the distance in $G$ between consecutive vertices of $P$ is at most $r$. For each $i\in[\ell]$, let $e_{i}=u_{i}u_{i+1}$ and $Q_{i}$ denote an $(u_{i},u_{i+1})$-isometric path in $G$. Finally, let $Q$ be the path in $G$ obtained by concatenating $Q_{1},Q_{2},\ldots,Q_{\ell}$. Observe that $\forall i\in[\ell],|E(Q_{i})|\leq r$ and therefore $|E(Q)|\leq\ell\cdot r$
[…]
Claim 5.14.
Every path $X\in\overline{\mathcal{Z}}$ is an isometric path in $G$.
Observe that, from the description of Procedure 2 and the definitions, any two paths in $\mathcal{Y}$ are edge disjoint. Hence, the paths in $\mathcal{Y}$ can be ordered as $Y_{1},Y_{2},\ldots,Y_{p}$ such that for $1\leq I<j\leq p$ all edges of $Y_{i}$ appear before those of $Y_{j}$ while traversing from one end-vertex of $Q$ to the other. Without loss of generality, we assume that one end-vertex of $Y_{1}$ is $u_{1}$. Hence, one end-vertex of $Y_{p}$ must be $u_{\ell+1}$. Let the two end-vertices of $Y_{i}$ be denoted as $a_{i}$ and $b_{i}$. Note that $a_{i+1}=b_{i}$ and without loss of generality we can assume that $a_{1}=u_{1}$, and $b_{p}=u_{\ell+1}$. Note that $\sum\limits_{i=1}^{p}|E(Y_{i})|=|E(Q)|\leq\ell\cdot r$. Using the above notations, we prove the following claim.
Claim 5.15.
$|\mathcal{Y}|\leq 2r-1.$
Proof.
Observe that $\left|\overline{\mathcal{Z}}\right|\leq\left|\mathcal{Z}\right|+1$. Hence proving $|\mathcal{Z}|\leq r-1$ is enough to prove 5.15. Assume for contradiction that $|\mathcal{Z}|\geq k$. Let $X_{i}$ denote an $(a_{i},b_{i})$-isometric path in $G$ and let $Q^{\prime}$ be the path between $u_{1}$ and $u_{\ell+1}$ obtained by concatenating $X_{1},X_{2},\ldots,X_{p}$. Observe that if for some $i\in[p]$, $Y_{i}\in\overline{\mathcal{Z}}$, then due to 5.14, $|E(X_{i})|=|E(Y_{i})|$. On the other hand, if for some $i\in[p]$, $Y_{i}\in\mathcal{Z}$, then $|E(X_{i})|\leq|E(Y_{i})|-1$. Hence,
$|E(Q^{\prime})|$ $=\displaystyle\sum\limits_{Y_{i}\in\overline{\mathcal{Z}}}|E(X_{i})|+\displaystyle\sum\limits_{Y_{i}\in\mathcal{Z}}|E(X_{i})|$
$\leq\displaystyle\sum\limits_{Y_{i}\in\overline{\mathcal{Z}}}|E(Y_{i})|+\displaystyle\sum\limits_{Y_{i}\in\mathcal{Z}}\left(|E(Y_{i})|-1\right)$
$\leq\displaystyle\sum\limits_{Y_{i}\in\overline{\mathcal{Z}}}|E(Y_{i})|+\displaystyle\sum\limits_{Y_{i}\in\mathcal{Z}}|E(Y_{i})|-k$
$\leq\displaystyle\sum\limits_{i}^{p}|E(Y_{i})|-k$
$\leq r(\ell-1)$
Hence $Q^{\prime}$ is a path between $u_{1}$ and $u_{\ell+1}$ of length at most $r(\ell-1)$. Hence $\mathsf{d}\left(u_{1},u_{\ell+1}\right)$ is at most $r(\ell-1)$. However, this contradicts 1. ∎
Now for each $Y_{i}\in\overline{\mathcal{Z}}$, define $Y^{\prime}_{i}=Y_{i}$ and for each $Y_{i}\in\mathcal{Z}$, define $Y^{\prime}_{i}=Y_{i}\setminus\{b_{i}\}$. Observe that each path in $\mathcal{Y}^{\prime}=\{Y^{\prime}_{i}\colon Y_{i}\in\mathcal{Z}\}\cup\{Y^{\prime}_{i}\colon Y_{i}\in\overline{\mathcal{Z}}\}$ is an isometric path in $G$. Due to 5.15, $|\mathcal{Y}^{\prime}|\leq 2r-1$. Also, $V(P)\subseteq V(Q)\subseteq\displaystyle\bigcup\limits_{Y\in\mathcal{Y}^{\prime}}V(Y)$. Therefore, the vertices of $P$ can be covered with $12k-$ 1 many isometric paths in $G$. ∎
Completion of the proof of Theorem 4(a): Let $G$ be a graph with $ipco\left(G\right)\leq k$ and $H$ be the $r^{th}$ power of $G$. Let $v$ be a vertex of $G$ such that the vertices of any isometric path of $G$ can be covered with $k$ many $v$-rooted isometric paths. Let $P$ be an isometric path of $H$. Due to Lemma 7 $P$ contains at most one black edge $e$. Let $P_{1}$ and $P_{2}$ be the two connected components obtained by removing $e$ from $P$. Note that both $P_{1}$ and $P_{2}$ are red isometric paths. Due to Lemma 8, for each $i\in\{1,2\}$, the vertices of $P_{i}$ can be covered with $12r-$ 1 many $v$-rooted isometric paths of $G$. Hence, the vertices of $P$ can be covered by $24r-$ 2 $v$-rooted isometric paths of $G$. This implies that the vertices of $P$ can be covered with $k(4r-2)$ many $v$-rooted isometric paths of $G$. Thus, Lemma 6 implies that $P$ can be covered by $k(4r^{2}-2r)$ $v$-rooted isometric paths of $H$. Hence, $ipco\left(H\right)\leq k(4r^{2}-2r)$. Similar arguments as above imply that if $sipco\left(G\right)\leq k$, then $sipco\left(H\right)\leq k{(4r^{2}-2r)}$.
4.2 Line graphs
In this section, we show that the (strong) isometric path complexity of the line graph of a graph $G$ is upper bounded by a constant factor of the (strong) isometric path complexity of $G$. First, we prove some lemmas.
Lemma 9.
Let $G$ be a graph and $P=e_{1}e_{2}\ldots e_{k}$ be an isometric path in $L\left(G\right)$. Then $e_{2},e_{3},\ldots,e_{k-1}$ induces an isometric path in $G$.
Proof.
Let $Q$ be the path induced by the edges $e_{2},e_{3},\ldots,e_{k-1}$ in $G$ and let $u,v$ be the end-vertices of $Q$. If $Q$ is not an isometric path, then another $(u,v)$-path $Q^{\prime}$ in $G$ of length at most $k-3$ exists. Let $e^{\prime}_{1},\ldots,e^{\prime}_{t}$ with $t\leq k-3$ be the edges of $Q^{\prime}$ and notice that $u$ (resp. $v$) is an end-vertex of $e^{\prime}_{1}$ (resp. $e^{\prime}_{k-3}$). Then $e_{2}e^{\prime}_{1}\ldots e^{\prime}_{t}e^{\prime}_{k-1}$ is a shorter path in $L\left(G\right)$, a contradiction. ∎
Lemma 10.
Let $G$ be a graph and $r$ be a vertex such that $ipco\left(\overrightarrow{G_{r}}\right)=k$. For any isometric path $P$, there exists a set $\mathcal{P}$ of $r$-rooted isometric paths with $|\mathcal{P}|\leq k$ such that $|E(P)\setminus\displaystyle\bigcup\limits_{Q\in\mathcal{P}}E(Q)|\leq k$.
Proof.
Since $ipco\left(\overrightarrow{G_{r}}\right)=k$, (by definition) the vertices of $P$ can be covered by $k$ many $r$-rooted isometric paths. Let $\mathcal{P}$ be a set of $r$-rooted isometric paths with $|\mathcal{P}|\leq k$ such that $V(P)\subset\displaystyle\bigcup\limits_{Q\in\mathcal{P}}V(Q)$ and $|E(P)\setminus\displaystyle\bigcup\limits_{Q\in\mathcal{P}}E(Q)|$ is minimized. In other words, $\mathcal{P}$ is a set of $k$ many $r$-rooted isometric paths that cover all vertices of $P$ and cover as many edges of $P$ as possible. We shall show that $\mathcal{P}$ satisfies the lemma statement. Let $X=E(P)\setminus\displaystyle\bigcup\limits_{Q\in\mathcal{P}}E(Q)$ and assume for contradiction that $|X|\geq k+1$. Let the edges of $X$ be ordered as $e_{1},e_{2},\ldots,e_{|X|}$ as they are encountered while traversing $P$ from one end-vertex $a$ to the other end-vertex $b$. For $i\in[|X|]$, let $e_{i}=u_{i}v_{i}$ and while traversing $P$ from $a$ to $b$, we encounter the following ordering of the end-vertices of the edges in $X$: $u_{1},v_{1},u_{2},v_{2},\ldots,u_{|X|},v_{|X|}$. Let $\{i,j\}\subset[|X|]$ and $Q_{i}\in\mathcal{P}$ be a path that contains $u_{i}$. Observe that $Q_{i}$ does not contain $u_{j}$. Otherwise, we could augment $Q_{j}$ to get a set of $r$-rooted isometric paths of cardinality $k$ (different from $\mathcal{P}$), that covers more edges than $\mathcal{P}$. This implies $|\mathcal{P}|\geq k+1$, which is a contradiction. ∎
Lemma 11.
Let $P$ be an isometric path in $G$. Then $L\left(P\right)$ is an isometric path in $L\left(G\right)$.
Proof.
Let the edges of $P$ be $e_{1},e_{2},\ldots,e_{k}$ and the two end-vertices of $P$ are $u,v$. Then $L\left(P\right)$ has length $k-1$ in $L\left(G\right)$. If $L\left(P\right)$ is not an isometric path, then there exists a $(e_{1},e_{k})$-path in $L\left(G\right)$ that can be written as $e_{1}=e^{\prime}_{1}e^{\prime}_{2}\ldots e^{\prime}_{t}=e_{k}$ such that $t\leq k-2$. Then by Lemma 9, we have that the edges $e^{\prime}_{2},\ldots,e^{\prime}_{t-1}$ induce an isometric path $Q$ in $G$ of length $k-4$. Now, including $e_{1}$ and $e_{k}$ creates a $(u,v)$-path of length at most $k-2$, a contradiction. ∎
Lemma 12.
Let $G$ be a graph, $P$ be an $(u,v)$-isometric path in $G$ and $e_{0}$ be any edge incident to $u$ in $G$. Then, the vertices of $L\left(P\right)$ can be covered by two $e_{0}$-rooted isometric paths in $L\left(G\right)$.
Proof.
Let the edges of $P$ be ordered as $e_{1},e_{2},\ldots,e_{k}$ as they are traversed starting from $u$. We also assume that the end-vertices of $e_{i}$ are $u_{i}$ and $u_{i+1}$ for $i\in[k]$ where $u$ is closer to $u_{i}$ than $u_{i+1}$. Since $L\left(P\right)$ is an isometric path in $L\left(G\right)$ (due to Lemma 11), if $e_{1}=e_{0}$, then the statement is true. Now assume $e_{1}\neq e_{0}$. observe that if $e_{0}$ shares an end-vertex with any edge $e_{j}$ with $j\geq 2$, then $P$ is not an isometric path in $G$. Now, assume that $i$ is the minimum index such that the path induced by $\{e_{0},e_{1},\ldots e_{i}\}$ is not an isometric path in $L\left(G\right)$. From previous arguments, it follows that $i\geq 3$.
Let $Q=e_{0}e^{\prime}_{1}\ldots e^{\prime}_{t}e_{i}$ be an $(e_{0},e_{i})$-isometric path in $L\left(G\right)$. Observe that $t\leq i-2$ and by Lemma 9, the path induced in $G$ by the end-vertices of the edges $e^{\prime}_{1},\ldots,e^{\prime}_{t}$ is an isometric path $P^{\prime}$ in $G$. If the common vertex of $e^{\prime}_{t}$ and $e_{i}$ is $u_{i+1}$, then by appending $u$ with $P^{\prime}$, it is possible to construct an $(u,u_{i+1})$-isometric path of length at most $i-1$ in $G$, which contradicts that $P$ is an isometric path. Hence, the common vertex of $e^{\prime}_{t}$ and $e_{i}$ is $u_{i}$. Let $Q^{\prime}$ denote the path $e_{0}e^{\prime}_{1}\ldots e^{\prime}_{t}e_{i}e_{i+1}\ldots e_{k}$ in $L\left(G\right)$ and observe that the length of $Q^{\prime}$ is at most $k-1$.
We argue that $Q^{\prime}$ is an isometric path in $L\left(G\right)$. If not, then there exists another isometric path $Q^{\prime\prime}=e_{0}e^{\prime\prime}_{1}\ldots e^{\prime\prime}_{x}e_{k}$ in $L\left(G\right)$ such that $x\leq k-3$. Due to Lemma 9, we know that the end-vertices of the edges $e^{\prime\prime}_{1},\ldots,e^{\prime\prime}_{x}$ induce an isometric path $P^{\prime\prime}$ in $G$. By including $u$ and $u_{k+1}$ to $P^{\prime\prime}$ it is possible to construct a $(u,u_{k+1})$-path in $G$ whose length is at most $k-1$. This is a contradiction. ∎
[…]
(a) (b)
Figure 6: (a) The graph $F_{5}$; (b) the graph $L\left(F_{5}\right)$.
Lemma 13.
Let $G$ be a graph with $ipco\left(G\right)=k$. Then $ipco\left(L\left(G\right)\right)\leq 3k+2$.
Proof.
Let $r$ be a vertex of $G$ such that $ipco\left(\overrightarrow{G_{r}}\right)=ipco\left(G\right)=k$. Let $e$ be any edge incident to $r$ and $Q=e_{1}e_{2}\ldots e_{x}$ be an isometric path in $L\left(G\right)$. Due to Lemma 9, $e_{2},e_{3},\ldots e_{x-1}$ induces an isometric path $P$ in $G$. Due to Lemma 10, there exists a set $\mathcal{P}$ of $r$ rooted isometric paths such that $|\mathcal{P}|\leq k$ and $|E(P)\setminus\displaystyle\bigcup\limits_{P^{\prime}\in\mathcal{P}}E(P^{\prime})|\leq k$. Let $e^{\prime}_{1},e^{\prime}_{2},\ldots e^{\prime}_{t}$ with $t\leq k$ are the edges of $P$ that do not belong to any path in $\mathcal{P}$. Due to Lemma 11, for each $P^{\prime}\in\mathcal{P}$, $L\left(P^{\prime}\right)$ is an isometric path in $L\left(G\right)$ and it can be decomposed into two $e$-rooted isometric paths, say $X(P^{\prime})$ and $Y(P^{\prime})$, in $L\left(G\right)$. Now consider the set $\mathcal{Q}=\displaystyle\bigcup\limits_{P^{\prime}\in\mathcal{P}}X(P^{\prime})\cup Y(P^{\prime})$. Notice that, it is possible that none of $E=\{e_{1},e^{\prime}_{1},\ldots,e^{\prime}_{t},e_{x}\}$ may lie in any path of $\mathcal{Q}$. For each $f\in E$, let $Z(f)$ denote an $(e,f)$-isometric path in $L\left(G\right)$. Clearly the set $\mathcal{Q}^{\prime}=\mathcal{Q}\displaystyle\bigcup\limits_{f\in E}Z(f)$ contains only $e$-rooted isometric paths, $|\mathcal{Q}^{\prime}|\leq 3k+2$ and any vertex of $Q$ lies in some isometric path in $\mathcal{Q}^{\prime}$. ∎
Similar arguments as above imply the following.
Lemma 14.
Let $G$ be a graph with $sipco\left(G\right)=k$. Then $sipco\left(L\left(G\right)\right)\leq 3k+2$.
Lemma 15.
For an integer $t\geq 1$, there exists a graph $G_{t}$ such that $ipco\left(G_{t}\right)=t$ and $ipco\left(L\left(G_{t}\right)\right)=2t-1$.
Proof.
For $t\geq 1$, the graph $F_{t}$ is defined in three steps as follows: (A) consider a path $P_{t}$ of length $12t-$ 1 and name the vertices $b_{1},c_{1},\ldots,b_{t},c_{t}$ where $b_{1}$ and $c_{t}$ are the two end-vertices of $P_{t}$; (B) consider a new vertex $r$ and for each $1\leq i\leq t$, introduce a new path $Q_{i}$ of length $t-1$ between $r$ and $b_{i}$; (C) introduce a new vertex $r_{0}$ adjacent to $r$. There are no other edges in $F_{t}$ except the ones mentioned above. Sometimes we shall call the paths $Q_{1},Q_{2},\ldots,Q_{t}$ pillars. Now construct $G_{t}$ by considering two disjoint copies (say $C^{1}$ and $C^{2}$) of $F_{t}$ and identify the vertices with label $r_{0}$.
We claim that $ipco\left(G_{t}\right)$ is $t$. Let $a\in V(G)$ be any vertex of $G_{t}$, and without loss of generality, assume that $a\in V(C^{1})$. Consider the copy of the path $P_{t}$ in $C^{2}$ and call it $S$. Observe that $S$ is an isometric path in $G_{t}$ and for any $\{i,j\}\subseteq[t]$, $\mathsf{d}\left(a,b_{i}\right)=\mathsf{d}\left(a,b_{j}\right)$. Moreover, for any $\{i,j\}\subseteq[t]$ there is no $(a,b_{i})$-isometric path in $G_{t}$ that contains $b_{j}$. Therefore $ipco\left(G_{t}\right)\geq t$.
Conversely, for $i\in\{1,2\}$, let $v^{i}$ denote the copy of the vertex $v\in V(F_{t})$ in $C^{i}$. (Recall that $C^{i}$ is isomorphic to $F_{t}$.) Let $T$ be any isometric path in $G_{t}$. If $T$ does not contain the vertex $r_{0}$, it must lie completely in $C^{1}$ or completely in $C^{2}$. Without loss of generality, assume $T$ lies in $C^{1}$. Then the vertices of $T$ can be covered by the set: $\{(r_{0},c^{1}_{j})-\text{isometric paths}\colon j\in[t]\}$. Now consider the case when $T$ contains $r_{0}$. Let $T^{i}=V(T)\cap V(C^{i})$. Observe that $T^{i}$ does not contain vertices from two different pillars. In this case, all the vertices of $T^{i}$ can be covered with an $(r_{0},c^{1}_{j})$-isometric path for some $j\in[t]$. Hence, the vertices of $T$ can be covered with two $r_{0}$-rooted isometric paths in $G_{t}$. Hence $ipco\left(G_{t}\right)\leq t$.
Now we claim that $ipco\left(L\left(G_{t}\right)\right)$ is $12t-$ 1. Observe that $L\left(G_{t}\right)$ is constructed by taking $L\left(C^{1}\right)$ and $L\left(C^{2}\right)$ and identifying the vertices corresponding to the edges $e_{0}=rr_{0}$. For $i\in\{1,2\}$, $L\left(C^{i}\right)$ can be described as follows.
  1. (a)
    Consider a clique $X$ of size $t+1$ whose vertices correspond to the edges $rr_{0},\{ra_{j}\}_{j\in[t]}$. Consider the pillar $Q_{j}$ and $a^{\prime}_{j}$ be the vertex adjacent to $a_{j}$ in $Q_{j}$. Observe that $ra_{j}$ is adjacent to $a_{j}a^{\prime}_{j}$ in $L\left(C^{i}\right)$. Let $H$ denote the union of $X$ and the edges between $ra_{j}$ and $a_{j}a^{\prime}_{j}$ in $L\left(C^{i}\right)$.
  2. (b)
    For each $j\in[t]$, let $b^{\prime}_{j}$ denote the vertex adjacent to $b_{j}$ in $Q_{j}$ and $f_{j}$ denote the edge $b_{j}b^{\prime}_{j}$. Let $e_{1}$ denote the edge $b_{1}c_{1}$, and for $j\in[2,t]$, let $e_{j},e_{j^{\prime}}$ denote the edges $b_{j}c_{j-1}$ and $b_{j}c_{j}$, respectively. Observe that $L\left(P_{t}\right)$ consists of the vertices $e_{1},e_{2},e^{\prime}_{2},\ldots,e_{t},e^{\prime}_{t}$. Moreover, $f_{1}$ is adjacent to $e_{1}$ in $L\left(G_{t}\right)$ and for each $j\in[2,t]$, $f_{j}$ is adjacent to both $e_{j},e^{\prime}_{j}$ in $L\left(G_{t}\right)$.
  3. (c)
    Finally, for each $j\in[t]$, consider the path $Q_{j}$. Then $L\left(C^{i}\right)$ is constructed by taking the union of $H,\left\{L\left(Q_{j}\right)\colon j\in[t]\right\}$ and $L\left(P_{t}\right)$.
We prove the following claims.
Claim 5.16.
For $i\in\{1,2\}$, the copy of $L\left(P_{t}\right)$ in $L\left(C^{i}\right)$ (say $S^{i})$ is an isometric path in $L\left(G_{t}\right)$. Moreover, the vertices $S^{i}$ are equidistant from the vertex $e_{0}$ in $L\left(G_{t}\right)$.
Proof.
Fix an $i\in\{1,2\}$. Since the copy of $P_{t}$ in $C^{i}$ is an isometric path in $G_{t}$, due to Lemma 11, $S_{i}$ is an isometric path in $L\left(G_{t}\right)$. Consider a pillar $Q_{j}$, $j\in[t]$, in $C^{i}$. Observe that $f_{j}$ (which corresponds to the edge $b_{j}b^{\prime}_{j}$ in $G_{t}$) is at distance $t$ from $e_{0}$ in $L\left(G_{t}\right)$. For $j=1$, $e_{j}$ is at distance $t+1$ from $e_{0}$ and for $j\in[2,t]$, both $e_{j},e^{\prime}_{j}$ are at distance $t+1$ from $e_{0}$. ∎
Using the above claim, we complete the proof. Let $e\in V(L\left(G_{t}\right))$ and without loss of generality assume $e\in V(L\left(C_{1}\right))$. Let $S$ denote the copy of the path $L\left(P_{t}\right)$ in $L\left(C_{2}\right)$. Due to 5.16, $S$ is an isometric path in $L\left(G_{t}\right)$ and all vertices of $L\left(P_{t}\right)$ are equidistant from the vertex $e_{0}=rr_{0}$ in $L\left(G_{t}\right)$. Moreover, there is no isometric path in $L\left(G_{t}\right)$ that contains two vertices of $S$. Therefore, at least $12t-$ 1 many $e$-rooted isometric paths are required to cover the vertices of $S$. Hence $ipco\left(L\left(G_{t}\right)\right)\geq 2t-1$.
Conversely, let $T$ be an isometric path in $L\left(G_{t}\right)$. If $T$ does not contain the vertex $e_{0}$, then the vertices of $T$ lie completely in $L\left(C_{1}\right)$ or completely in $L\left(C_{2}\right)$. In either case, the vertices of $T$ can be covered with $12t-$ 1 many $e_{0}$-rooted isometric paths. If $T$ contains the vertex $e_{0}$, then $T$ can be covered by two $e_{0}$-rooted isometric paths in $L\left(G_{t}\right)$. Hence $ipco\left(L\left(G_{t}\right)\right)\leq 2t-1$. ∎
The proof of Theorem 4(b) follows from Lemmas 13, 14 and 15.
4.3 Clique-sum
In this section, we prove Theorem 5. We begin by proving the following lemma.
Lemma 16.
Let $H$ be the clique-sum of two graphs, $H_{1}$ and $H_{2}$, with $C$ being the common clique. Let $P$ be an isometric path of $H$ such that $V(P)\subseteq V(H_{1})$, $P$ has an end-vertex in $C$, and $w$ be any vertex of $H_{2}$. Then, the vertices of $P$ can be covered with three $w$-rooted isometric paths in $H$.
Proof.
Assume for contradiction that there are no three $w$-rooted isometric paths in $H$ that cover all vertices of $V(P)$. Let $u$ and $v$ be the end-vertices of $P$ and $v\in C$. For some isometric path $T$ and two vertices $\{x,y\}\subseteq V(T)$, $T\left(x,y\right)$ shall denote the subpath of $T$ between $x$ and $y$. Let $u_{1}=u$, and for $i\in\{1,2,3\}$, let $Q_{i}$ be a minimal $w$-rooted isometric path that contains $u_{i}$ and covers the maximum number of vertices of $P$, and define $u_{i+1}$ to be the vertex in $V(P)\setminus\displaystyle\bigcup\limits_{j<i+1}V(Q_{j})$ that is at minimum distance from $u$. For $i\in\{1,2,3\}$, $v_{i}\in V(Q_{i})\cap C$ that is at minimum distance from $w$.
Claim 5.17.
The vertex $u$ is an end-vertex of $Q_{1}$.
Proof.
Observe that $V(Q_{1})\cap V(P)$ must be a subpath of $P$ and the minimality of $Q_{1}$ implies that the end-vertex of $Q_{1}$ must be in $V(Q_{1})\cap V(P)$. Let $x\neq w$ be an end-vertex of $Q_{1}$ and suppose $x\neq u$. Let $\mathsf{d}\left(v,x\right)=a$ and as $x\neq u$, $\mathsf{d}\left(v,u\right)\geq a+1$. Then $\mathsf{d}\left(v_{1},x\right)\leq a+1$ and furthermore $\mathsf{d}\left(v_{1},u\right)\leq a$. This implies $\mathsf{d}\left(v,u\right)=a+1,\mathsf{d}\left(v_{1},x\right)=a+1$. Now consider the path $T$ induced by $V(Q_{1}\left(w,v_{1}\right))\cup V(P\left(v,x\right))$. Clearly, $T$ is an $(w,x)$-isometric path ad together with $Q_{1}\left(w,u\right)$ covers all vertices of $P$, which contradicts our assumption that $P$ cannot be covered with three $w$-rooted paths. ∎
Similar arguments, as used in the proof of 5.17, imply the following.
Claim 5.18.
The vertices $u_{2}$ and $u_{3}$ are end-vertices of $Q_{2}$ and $Q_{3}$, respectively.
For $i\in\{1,2,3\}$, let $u^{\prime}_{i}$ be the vertex in $V(P)\cap V(Q_{i})$ which is adjacent to $u_{i+1}$. Let $\mathsf{d}\left(u^{\prime}_{1},v\right)=a$, $\mathsf{d}\left(w,v_{1}\right)=b$, and $\mathsf{d}\left(u^{\prime}_{2},u_{2}\right)=\ell$. 5.18 implies that the vertices $u_{1}=u,u^{\prime}_{1},u_{2},u^{\prime}_{2},u_{3},u^{\prime}_{3},u_{4}$ appear in that order when $P$ is traversed from $u$ to $v$.
Claim 5.19.
$\mathsf{d}\left(v_{1},u^{\prime}_{1}\right)=a-1$.
Proof.
Since $\mathsf{d}\left(v,u^{\prime}_{1}\right)=a$ and $v$ is adjacent to $v_{1}$, we have $a-1\leq\mathsf{d}\left(u^{\prime}_{1},v_{1}\right)\leq a+1$. If $\mathsf{d}\left(v_{1},u^{\prime}_{1}\right)=a+1$, then $V(Q_{1}\left(w,v_{1}\right))\cup V(P)$ induces a $(w,u_{1})$-isometric path that covers strictly more vertices of $P$ than $Q_{1}$, which is a contradiction. Now assume, $\mathsf{d}\left(v_{1},u^{\prime}_{1}\right)=a$. observe that the path given by $V(P\left(u_{2},v\right))\cup V(Q_{1}\left(v_{1},w\right))$ cannot be isometric (otherwise, two $w$-rooted isometric paths cover all vertices of $P$). Hence, $\mathsf{d}\left(u_{2},w\right)\leq a+c-1$ and let $Z$ be any $(w,u_{2})$-isometric path. observe that $Z\cup\left\{u^{\prime}_{1}\right\}$ is a $(w,u^{\prime}_{1})$-isometric path and therefore, $V(Z)\cup V(P\left(u_{2},u\right))$ is an $(w,u)$-isometric path that covers strictly more vertices of $P$ than $Q_{1}$, a contradiction. ∎
Claim 5.20.
$\mathsf{d}\left(u_{2},w\right)=a+b-1$.
Proof.
5.19 implies that $a+b-2\leq\mathsf{d}\left(w,u_{2}\right)\geq a+b$. If $\mathsf{d}\left(w,u_{2}\right)=a+b-2$, then any $(w,u_{2})$-isometric path concatenated with $P\left(u_{2},u\right)$ yields a $(w,u)$-isometric path that covers strictly more vertices of $P$ than $Q_{1}$, which is a contradiction. If $\mathsf{d}\left(u_{2},w\right)=a+b$, then $V(P\left(u_{2},v\right))\cup V(Q_{1}\left(v_{1},w\right))$ yields an isometric path that covers strictly more vertices of $P\left(u_{2},v\right)$ than $Q_{1}$, a contradiction. ∎
Recall that $\mathsf{d}\left(u^{\prime}_{2},u_{2}\right)=\ell$ and observe that $\mathsf{d}\left(v,u_{3}\right)=a-\ell-2$.
Claim 5.21.
$\mathsf{d}\left(u_{3},w\right)=a+b-\ell-1$.
Proof.
Observe that $\mathsf{d}\left(v,u_{3}\right)=a-1-\ell$. If $\mathsf{d}\left(u_{3},w\right)=a+b-\ell$, then $V(Q_{1}\left(v_{1},w\right))\cup V(P\left(v,u_{3}\right))$ yields an isometric path that covers strictly more vertices of $P\left(u_{3},v\right)$ than $Q_{3}$, a contradiction. If $\mathsf{d}\left(u_{3},w\right)=a+b-\ell-2$, then 5.20 implies that $V(Q_{3}\left(w,u_{3}\right))\cup V(P\left(u^{\prime}_{2},u_{2}\right))$ induces an $(w,u_{2})$-isometric path that covers strictly more vertices than $Q_{2}$, a contradiction. ∎
Now consider the path $T$ induced by $Q_{1}\left(w,v_{1}\right)\cup P\left(v,u_{3}\right)$. The length of $T$ is at most $b+a-\ell-1$ and therefore is an $(w,u_{3})$-isometric path that covers strictly more vertices than $Q_{3}$, a contradiction. ∎
Let $G_{1},G_{2},\ldots,G_{t}$ be graphs, each with strong isometric path complexity at most $k$ and $G=\displaystyle\bigoplus\limits_{i=1}^{t}G_{i}$. Let $P$ be any isometric path of $G$ and $v$ be any vertex of $G$. We shall show that the vertices of $P$ can be covered with at most $183k+$ 18 many $v$-rooted isometric paths. First, we make the following observations.
Observation 2.
For a subset $X\subseteq[t]$, $\bigoplus\limits_{i\in X}G_{i}$ is an isometric subgraph of $G$.
Observation 3.
For an integer $i\in[t]$, $V(P)\cap V(G_{i})$ is either empty or induces a subpath of $P$.
Lemma 17.
Let $w$ be a vertex such that there exists an integer $i\in[t]$ such that $V(P)\cap V(G_{i})$ is non-empty and $w\in V(G_{i})$. Then, the vertices of $P$ can be covered with $k+6$ many $w$-rooted isometric paths in $G$.
Proof.
3 implies that $V(P)\cap V(G_{i})$ induces a subpath, say $P_{i}$, of $P$. Let $Q$ be a connected component of $V(P)\setminus V(P_{i})$. observe that there exists a clique $C$ in $G$ that separates $Q$ from $V(G_{i})\setminus V(C)$. Let $H_{1}$ and $H_{2}$ be the two connected components in $V(G)\setminus C$ that contain $Q$ and $V(G_{i})\setminus V(C)$, respectively. Let $H$ be the clique-sum of $H_{1},H_{2}$ and $C$ be the common clique. By 2, $H$ is an isometric subgraph of $G$.
Observe that one end-vertex of $Q$ is adjacent to exactly one vertex of $C$, say $u$, of $P_{i}$. Let $Q^{+}$ denote the path induced by $V(Q)\cup\{u\}$. Observe that $Q^{+}$ is an isometric path of $H$, $V(Q^{+})\subseteq V(H_{1})$, and $w$ is a vertex of $H_{2}$. Now, Lemma 16 implies that the vertices of $Q^{+}$ can be covered with three $w$-rooted isometric paths of $H$.
To complete the proof of the lemma, observe that there can be at most two connected components of $V(P)\setminus V(G_{i})$. The above arguments imply that all vertices of $V(P)\setminus V(P_{i})$ can be covered with six $w$-rooted isometric paths in $G$. Since $sipco\left(G_{i}\right)\leq k$, the vertices of $P_{i}$ can be covered with at most $k$ many $w$-rooted isometric paths of $G$. Hence, $P$ can be covered with $k+6$ many $w$-rooted isometric paths of $G$. ∎
Completion of the proof of the theorem: Observe that if there exists an integer $i\in[t]$ such that $V(P)\cap V(G_{i})$ is non-empty and $v\in V(G_{i})$, then the theorem follows from Lemma 17. Otherwise, there exists an integer $i\in[t]$ such that $V(G_{i})\cap V(P)$ is non-empty and there exists a clique $C$ such that $v$ and $P$ lies in different connected components in the graph induced by $V(G)\setminus C$. Let $v^{\prime}$ be any vertex of $C$. Lemma 17 implies that there is a set $\mathcal{P}$ of $k+6$ many $v^{\prime}$-rooted isometric paths in $G$ such that all vertices of $P$ are covered by the paths in $\mathcal{P}$. Now, by Lemma 16, each path in $\mathcal{P}$ can be covered with three $v$-rooted isometric paths in $G$. Therefore, the vertices of $P$ can be covered with $183k+$ 18 many $v$-rooted isometric paths in $G$.
5 Conclusion
We proved that $U_{t}$-asymptotic minor-free graphs have bounded strong isometric path complexity. However, it is easy to construct a graph class whose strong isometric path complexity is small but contains all $U_{t}$-asymptotic minors. For $t\geq 1$, consider $K_{1,t}$ and replace its edges with a path of length $t$. Let $u_{1},u_{2},\ldots,u_{t}$ denote the leaves. Let $H_{t}$ denote the graph obtained by introducing a path of length $t$ between $u_{i}$ and $u_{i+1}$ for each $i\in[t-1]$. Observe that $sipco\left(H_{t}\right)\leq 2$, but =$\mathcal{H}=\{H_{t}\}_{t\in\mathbb{N}}$ contains $U_{t}$-asymptotic minors for all $t$. Therefore, $U_{t}$-asymptotic minors do not provide a structural characterization of graphs with bounded strong isometric path complexity. It would be interesting to obtain such a characterization.
We proved that important graph operations like fixed powers, line graphs, and clique-sum preserve strong isometric path complexity. However, it is not clear if our upper bounds are tight. It would also be interesting to characterize which graph operations preserve strong isometric path complexity.
Graphs with bounded strong isometric path complexity generalize both hyperbolic graphs and $U_{t}$-asymptotic minor-free graphs and their subclasses. They can be seen as an extension of hyperbolic graphs, where arbitrarily large isometric cycles are allowed to exist (although in a restricted fashion).
It is more generally of interest to see which interesting properties can be lifted from them upto graphs of bounded strong isometric path complexity. This could shed new light on these classes.
It would also be of interest to discover other algorithmic applications of (strong) isometric path complexity beyond Isometric Path Cover.